5. If the manager of a paint supply store wants to estimate, with 95% confidence, the mean amount of paint in a 1-gallon can to within ± 0.004 gallon and also assumes that the standard deviation is 0.02 gallon, what sample size is needed?
6. You want to rent an unfurnished one-bedroom apartment for next semester. The mean monthly rent for a random sample of 10 apartments advertised in the local newspaper is $540. Assume that the 휎=$80.a) Find a 95% confidence interval for the mean monthly rent for unfurnished one-bedroom apartments available for rent in this community.b) If we increase the sample size to 100 apartments, the sdandard error becomes ________, then with a fixed confidence level, the margin of error E becomes _______________ and finally the interval is ________________.
1
Expert's answer
2019-11-05T10:08:14-0500

5 . Sample = "n =(\frac {\sigma*z} {me})^2 = \frac{(0.02*1.96)^2}{0.004^2} = 96.04, \therefore n =97"


6 . (a) 95% confidence interval ="540± 1.96 \times \frac{80}{\sqrt{10}} = 540 ±49.58 = (490.42, 589.58)"

(b) standard error = "\frac{80} {\sqrt{100}}=8"

margin of error ="E=1.96\times\frac{80} {\sqrt{100}}=1.96\times8=15.68"

95% confidence interval ="540 \u00b1 15.68 = (524.32, 555.68)"